### v2

@@ Abstract
To construct responsible and secure AI applications, harmful information data is widely utilized for adversarial testing and the development of safeguards. Existing studies mainly leverage Large Language Models (LLMs) to synthesize data to obtain high-quality task datasets at scale, thereby avoiding costly human annotation. However, limited by the safety alignment mechanisms of LLMs, the synthesis of harmful data still faces challenges in generation reliability and content diversity. In this study, we propose a novel harmful information synthesis framework, PoisonSwarm, which applies the model crowdsourcing strategy to generate diverse harmful data while maintaining a high success rate. Specifically, we generate abundant benign data as the based templates in a counterfactual manner. Subsequently, we decompose each based template into multiple semantic units and perform unit-by-unit toxification and final refinement through dynamic model switching, thus ensuring the success of synthesis. Experimental results demonstrate that PoisonSwarm achieves state-of-the-art performance in synthesizing different categories of harmful data with high scalability and diversity. Warning: This paper has certain harmful content to serve as examples.
Keywords:
Harmful Information Detection Data Synthesis
@@ 1 Introduction
Building on the strong-weak model collaboration framework, we propose PoisonSwarm, a novel approach for robustly synthesizing diverse harmful data through model crowdsourcing. In PoisonSwarm, LLMs collaborate to generate abundant benign data and progressively toxify it, thus creating diverse harmful data while maintaining a high success rate. Specifically, we first employ counterfactual methods to generate benign content that has the desired structure and thematic backbone. This benign content is then decomposed into smaller semantic units, each of which is assigned to different LLMs for toxicification. If a unit produces invalid results or malfunctions, PoisonSwarm dynamically switches to another LLM to continue the toxification process. Finally, all toxicified content is integrated and refined into high-quality harmful data. Experimental results demonstrate that PoisonSwarm outperforms existing methods across multiple evaluation metrics, highlighting its effectiveness in generating diverse and reliable harmful data. Our major contributions are:
@@
  • This study proposes the universal harmful information synthesis approach PosionSwarm, which generates abundant benign data as the based templates and toxicifies them via crowdsourcing them to multiple LLMs, thus handling the challenge of content diversity and generation reliability.
  • Experimental results demonstrate that PoisonSwarm outperforms existing methods in harmful information synthesis across multiple evaluation metrics, highlighting its effectiveness in generating diverse and high-quality harmful data for adversarial training and testing, thereby enhancing AI security.
@@ 3 Motivation
We focus on hate speech for illustration, which is the typical harmful information. As shown in Fig.LABEL:Fig.mot, we systematically compare the effectiveness of existing harmful content detectors in identifying human-generated and AI-generated harmful information. Specifically, for hate-speech data, we utilize PoisonSwarm to generate 90 instances of hate-speech targeting a simulated user Johnny. Table 1 presents representative examples of synthesized data. Additionally, we further generate 40 counter-narrative and supportive speeches to test the systems’ false positive rates on AI-generated harmful content. The human-generated dataset is sourced from Measuring Hate Speech (MHS) [9], providing a benchmark for comparison. For detectors, we select models from huggingface, i.e., GroNLP/hateBERT, MilaNLProc/hate-ita, Hate-speech-CNERG/bert-base-uncased-hatexpla
@@
The detection results presented in Fig.LABEL:Fig.mot demonstrate that generative AI can generate more implicit and informationally rich harmful content, as shown in cloud word, which can effectively evade the identification of traditional detectors that rely on fixed linguistic patterns such as toxic keywords (e.g., "f**k", "s**t"). Consequently, these conventional detection methods, which are primarily trained on limited harmful information with explicit lexical features, exhibit diminished performance when confronted with AI-generated harmful content. To alleviate this vulnerability, it is important to develop methods for obtaining more data that comprehensively reflects the diverse patterns of harmful information, thereby providing fully adversarial training and testing for AI safeguards.
@@ 4 Methodology
In this section, we introduce our PoisonSwarm, which can synthesize diverse harmful information while maintaining a high success rate by using strong-weak model collaboration. The overall structure of PoisonSwarm is shown in Fig.3.
Figure 3: The overall structure of our PoisonSwarm. Specifically, Model Crowdsourcing Queue (§4.1) is the core mechanism of PoisonSwarm to ensure the success of obtaining harmful content. In Counterfactual Mapping (§4.2), PoisonSwarm first generates benign content with the desired structure and thematic elements. Then, in Adversarial Content Toxicifying (§4.3), the benign content is segmented into smaller semantic units for targeted toxification. Multiple LLMs collaborate to transform these units into harmful content, with dynamic model switching when malfunctions occur. Finally, in Hallucination Evaluation (§4.4), the generated harmful information is validated by its coherence and relevance.
@@ 4.1 Model Crowdsourcing Queue
  • Fallback Models (FMs, $\mathcal{M}_{F}$) are LLMs that reliably address the harmful queries in PoisonSwarm even in greedy decoding (temperature is 0). FMs serve as fallback options when higher-level models fail to produce acceptable outputs, including Qwen{2,2.5}-7B-Instruct and GLM4-9B-chat.
@@ 4.2 Counterfactual Mapping
While the Model Crowdsourcing strategy leverages vulnerabilities in different LLMs to address harmful queries, it still cannot guarantee that our PoisonSwarm can benefit from advanced aligned LLMs. Therefore, we employ Counterfactual Mapping (CM) to decouple harmless subtasks from the harmful data synthesis process, enabling these advanced LLMs to contribute effectively.
@@ 5.1.2 Baseline Methods
We compare our PoisonSwarm with existing data augmentation and synthesis methods. We use typical data augmentation methods from nlpaug [11], including Synonym Augmentation (Syn.Aug.), Contextual Word Embeddings Augmenter (Token.Aug.), and Contextual Word Embeddings for Sentence Augmenter (Sen.Aug.). In terms of data synthesis methods, we directly query the LLM to synthesize toxic content as the baseline, i.e., DQ-LLM. Regarding the bypass of the safety alignment mechanisms of LLMs, we compare PoisonSwarm with black-box LLM attacking methods as the strong baselines, which are used to jailbreak LLMs for harmful data synthesis: PAIR [3] generates semantic prompt-level jailbreaks with an attacker LLM. TAP [12] employs tree-structured prompting to elicit harmful behaviors. LLM-Fuzzer [17] automates the generation of jailbreak prompts using a black-box fuzzing framework, iteratively mutating human-written templates to attack LLMs.
@@ 5.1.4 Experimental Setups
The embedder we employed is bert-base-uncased [5]. PoisonSwarm adopts gpt-4o, gpt-4o-mini, and qwen2.5-7B-Instruct as crowdsourced models, while other LLM-driven baselines adopt gpt-4o-mini. We construct the data synthesis prompts using few-shot in-context learning, where each prompt targets one specific subgroup as defined in MHS. We calculate Div. by computing the cosine similarity based on text embeddings from BGE-M3 [4]. We calculate Tox. and Nat. using gpt-4o. In the evaluating stage (metric calculation, hallucination evaluation), the temperature of LLM is 0.0, otherwise, it is 0.7.
@@ 5.2 Experimental Results
To comprehensively evaluate PoisonSwarm’s capabilities, our experiments are designed to address these research questions (RQs):
  • RQ1: Does PoisonSwarm outperform existing methods? What mechanisms contribute to its success?
@@
  • RQ3: How does PoisonSwarm ensure the diversity of synthesized harmful information?
Baseline Comparison (RQ1.1) To investigate RQ1, we first conduct comprehensive experiments with different baselines and variants of PoisonSwarm for comparison. We use different methods to generate 100 hateful samples for each of the target categories (4,600 samples in total), aiming to validate their effectiveness in generating harmful content.
@@
As shown in Table 3, we observe different trade-off patterns between data augmentation and data synthesis methods. The data generated from traditional data augmentation methods (e.g., Syn.Aug. and Sen.Aug.) achieves superior toxicity scores (0.7627 and 0.7660), while it is highly different from human expression with low naturalness (0.2393 and 0.1275). In contrast, the data synthesized from jailbroken LLMs achieves superior naturalness by leveraging their pre-trained knowledge, but it struggles with maintaining harmful characteristics due to their inherent training goal of generating helpful and safe content. Even when jailbroken to bypass safety alignment mechanisms, the LLM still tends to be conservative in toxicity generation without explicit harmful references, as evidenced by the high naturalness and lower toxicity of DQ-LLM, PAIR, TAP, and LLM-Fuzzer. PoisonSwarm effectively achieves the best trade-off with balanced performance across all metrics. It achieves the highest success rate (0.9035) and diversity score (0.5606), while maintaining competitive toxicity (0.6860) and naturalness (0.6583), highlighting the design of harmful data generation by speculatively toxicifying the diverse harmless templates using model crowdsourcing.
Ablation Study (RQ1.2) Then, we further explore the contribution of each component in PoisonSwarm by ablation study, as shown in Table 4. The ablation settings and analysis are as follows:
@@
As shown in Fig.5, models trained solely on human-generated data perform well on human-generated harmful content but struggle with AI-generated samples. Incorporating AI-generated data into the training set significantly improves the detection of AI-generated content while maintaining comparable performances on human-generated content. This phenomenon indicates significant differences in the expression patterns of harmful content between humans and AI. As shown in Table 6, human-generated harmful content tends to employ direct insulting terms and simple sentence structures (e.g., "scum"), which are obvious markers of aggression and thus being easier for detectors to identify. In contrast, AI-generated harmful content often adopts a seemingly rational tone to conceal biases, making it more adept at evading detection. These differences further highlight the necessity of our proposed PoisonSwarm, which provides richer AI-generated data for harmful information detection. As for our PoisonSwarm, the pairwise similarity distribution in Fig.LABEL:Fig.dis indicates that its synthesized content is less similar, and the t-SNE visualization in Fig.LABEL:Fig.tsne indicates that its synthesized content is the semantic complement for human-generated data.
@@
Case Study (RQ3) PoisonSwarm ensures the diversity of synthesized harmful information by leveraging diverse benign templates as the basis for dynamic toxification. As illustrated in Fig.LABEL:Fig.goodbad, the benign content template (left) contains a wide range of neutral and positive semantic units, contributing to the foundational diversity of the final generated harmful information (right). PoisonSwarm introduces the counterfactual mapping module to obtain diverse templates from strong LLMs without being constrained by LLM’s safe alignment mechanism. Then, PoisonSwarm transforms these benign templates into harmful expressions using dynamic model switching, thereby preserving semantic richness and avoiding homogeneity. Consequently, the resulting harmful content maintains high diversity, effectively reflecting a broad spectrum of harmful expressions.
@@ 6 Conclusion
In this work, we propose PoisonSwarm, a universal harmful information synthesis method utilizing model crowdsourcing to generate diverse harmful data while maintaining a high success rate. To benefit from the rapid development of strong LLMs for high-quality harmful information synthesis, we introduce the strong-weak model collaboration framework, where we decompose the harmful data synthesis tasks into benign template generation by strong LLMs and template toxicification by weak LLMs. We highlight significant differences between AI-generated and human-generated harmful information, which indicate the limitations of previous passive harmful data collection for detector construction and emerging threats from the misuse of generative AI for AI-driven harmful speech campaigns. Experimental evaluations clearly illustrate that PoisonSwarm achieves superior performance compared to existing harmful information generation methods, demonstrating its efficacy in synthesizing scalable and diverse harmful information for secure and responsible AI development.

### v3

@@ Abstract
To construct responsible and secure AI applications, harmful information data is widely utilized for adversarial testing and the development of safeguards. Existing studies mainly leverage Large Language Models (LLMs) to synthesize data to obtain high-quality task datasets at scale, thereby avoiding costly human annotation. However, limited by the safety alignment mechanisms of LLMs, the synthesis of harmful data still faces challenges in generation reliability and content diversity. In this study, we propose a novel harmful information synthesis framework, SwarmLaunder, which applies the model crowdsourcing strategy to generate diverse harmful data while maintaining a high success rate. Specifically, we generate abundant benign data as the based templates in a counterfactual manner. Subsequently, we decompose each based template into multiple semantic units and perform unit-by-unit toxification and final refinement through dynamic model switching, thus ensuring the success of synthesis. Experimental results demonstrate that SwarmLaunder achieves state-of-the-art performance in synthesizing different categories of harmful data with high scalability and diversity. Warning: This paper has certain harmful content to serve as examples.
@@ 1 Introduction
Building on the strong-weak model collaboration framework, we propose SwarmLaunder, a novel approach for robustly synthesizing diverse harmful data through model crowdsourcing. In SwarmLaunder, LLMs collaborate to generate abundant benign data and progressively toxify it, thus creating diverse harmful data while maintaining a high success rate. Specifically, we first employ counterfactual methods to generate benign content that has the desired structure and thematic backbone. This benign content is then decomposed into smaller semantic units, each of which is assigned to different LLMs for toxicification. If a unit produces invalid results or malfunctions, SwarmLaunder dynamically switches to another LLM to continue the toxification process. Finally, all toxicified content is integrated and refined into high-quality harmful data. Experimental results demonstrate that SwarmLaunder outperforms existing methods across multiple evaluation metrics, highlighting its effectiveness in generating diverse and reliable harmful data. Our major contributions are:
@@
  • This study proposes the universal harmful information synthesis approach SwarmLaunder, which generates abundant benign data as the based templates and toxicifies them via crowdsourcing them to multiple LLMs, thus handling the challenge of content diversity and generation reliability.
  • Experimental results demonstrate that SwarmLaunder outperforms existing methods in harmful information synthesis across multiple evaluation metrics, highlighting its effectiveness in generating diverse and high-quality harmful data for adversarial training and testing, thereby enhancing AI security.
@@ 3 Motivation
We focus on hate speech for illustration, which is the typical harmful information. As shown in Fig.3(a), we systematically compare the effectiveness of existing harmful content detectors in identifying human-generated and AI-generated harmful information. Specifically, for hate-speech data, we utilize SwarmLaunder to generate 90 instances of hate-speech targeting a simulated user Johnny. Table 1 presents representative examples of synthesized data. Additionally, we further generate 40 counter-narrative and supportive speeches to test the systems’ false positive rates on AI-generated harmful content. The human-generated dataset is sourced from Measuring Hate Speech (MHS) [9], providing a benchmark for comparison. For detectors, we select models from huggingface, i.e., GroNLP/hateBERT, MilaNLProc/hate-ita, Hate-speech-CNERG/bert-base-uncased-hatexpla
@@
The detection results presented in Fig.3(a) demonstrate that generative AI can generate more implicit and informationally rich harmful content, as shown in cloud word, which can effectively evade the identification of traditional detectors that rely on fixed linguistic patterns such as toxic keywords (e.g., "f**k", "s**t"). Consequently, these conventional detection methods, which are primarily trained on limited harmful information with explicit lexical features, exhibit diminished performance when confronted with AI-generated harmful content. To alleviate this vulnerability, it is important to develop methods for obtaining more data that comprehensively reflects the diverse patterns of harmful information, thereby providing fully adversarial training and testing for AI safeguards.
(a) Comparison of existing harmful information detector performance on human/AI-generated content. Fig.LABEL:Fig.face demonstrates the cloud word of hate-speech targeting simulated user Johnny by AI generating. Fig.LABEL:Fig.radar illustrates the significant performance gap across multiple detectors, highlighting the challenges posed by AI-driven harmful speech campaigns.
@@ 4 Methodology
In this section, we introduce our SwarmLaunder, which can synthesize diverse harmful information while maintaining a high success rate by using strong-weak model collaboration. The overall structure of SwarmLaunder is shown in Fig.3.
Figure 3: The overall structure of our SwarmLaunder. Specifically, Model Crowdsourcing Queue (§4.1) is the core mechanism of SwarmLaunder to ensure the success of obtaining harmful content. In Counterfactual Mapping (§4.2), SwarmLaunder first generates benign content with the desired structure and thematic elements. Then, in Adversarial Content Toxicifying (§4.3), the benign content is segmented into smaller semantic units for targeted toxification. Multiple LLMs collaborate to transform these units into harmful content, with dynamic model switching when malfunctions occur. Finally, in Hallucination Evaluation (§4.4), the generated harmful information is validated by its coherence and relevance.
@@ 4.1 Model Crowdsourcing Queue
  • Fallback Models (FMs, $\mathcal{M}_{F}$) are LLMs that reliably address the harmful queries in SwarmLaunder even in greedy decoding (temperature is 0). FMs serve as fallback options when higher-level models fail to produce acceptable outputs, including Qwen{2,2.5}-7B-Instruct and GLM4-9B-chat.
@@ 4.2 Counterfactual Mapping
While the Model Crowdsourcing strategy leverages vulnerabilities in different LLMs to address harmful queries, it still cannot guarantee that our SwarmLaunder can benefit from advanced aligned LLMs. Therefore, we employ Counterfactual Mapping (CM) to decouple harmless subtasks from the harmful data synthesis process, enabling these advanced LLMs to contribute effectively.
@@ 5.1.2 Baseline Methods
We compare our SwarmLaunder with existing data augmentation and synthesis methods. We use typical data augmentation methods from nlpaug [11], including Synonym Augmentation (Syn.Aug.), Contextual Word Embeddings Augmenter (Token.Aug.), and Contextual Word Embeddings for Sentence Augmenter (Sen.Aug.). In terms of data synthesis methods, we directly query the LLM to synthesize toxic content as the baseline, i.e., DQ-LLM. Regarding the bypass of the safety alignment mechanisms of LLMs, we compare SwarmLaunder with black-box LLM attacking methods as the strong baselines, which are used to jailbreak LLMs for harmful data synthesis: PAIR [3] generates semantic prompt-level jailbreaks with an attacker LLM. TAP [12] employs tree-structured prompting to elicit harmful behaviors. LLM-Fuzzer [17] automates the generation of jailbreak prompts using a black-box fuzzing framework, iteratively mutating human-written templates to attack LLMs.
@@ 5.1.4 Experimental Setups
The embedder we employed is bert-base-uncased [5]. SwarmLaunder adopts gpt-4o, gpt-4o-mini, and qwen2.5-7B-Instruct as crowdsourced models, while other LLM-driven baselines adopt gpt-4o-mini. We construct the data synthesis prompts using few-shot in-context learning, where each prompt targets one specific subgroup as defined in MHS. We calculate Div. by computing the cosine similarity based on text embeddings from BGE-M3 [4]. We calculate Tox. and Nat. using gpt-4o. In the evaluating stage (metric calculation, hallucination evaluation), the temperature of LLM is 0.0, otherwise, it is 0.7.
@@ 5.2 Experimental Results
To comprehensively evaluate SwarmLaunder’s capabilities, our experiments are designed to address these research questions (RQs):
  • RQ1: Does SwarmLaunder outperform existing methods? What mechanisms contribute to its success?
@@
  • RQ3: How does SwarmLaunder ensure the diversity of synthesized harmful information?
Baseline Comparison (RQ1.1) To investigate RQ1, we first conduct comprehensive experiments with different baselines and variants of SwarmLaunder for comparison. We use different methods to generate 100 hateful samples for each of the target categories (4,600 samples in total), aiming to validate their effectiveness in generating harmful content.
@@
As shown in Table 3, we observe different trade-off patterns between data augmentation and data synthesis methods. The data generated from traditional data augmentation methods (e.g., Syn.Aug. and Sen.Aug.) achieves superior toxicity scores (0.7627 and 0.7660), while it is highly different from human expression with low naturalness (0.2393 and 0.1275). In contrast, the data synthesized from jailbroken LLMs achieves superior naturalness by leveraging their pre-trained knowledge, but it struggles with maintaining harmful characteristics due to their inherent training goal of generating helpful and safe content. Even when jailbroken to bypass safety alignment mechanisms, the LLM still tends to be conservative in toxicity generation without explicit harmful references, as evidenced by the high naturalness and lower toxicity of DQ-LLM, PAIR, TAP, and LLM-Fuzzer. SwarmLaunder effectively achieves the best trade-off with balanced performance across all metrics. It achieves the highest success rate (0.9035) and diversity score (0.5606), while maintaining competitive toxicity (0.6860) and naturalness (0.6583), highlighting the design of harmful data generation by speculatively toxicifying the diverse harmless templates using model crowdsourcing.
Ablation Study (RQ1.2) Then, we further explore the contribution of each component in SwarmLaunder by ablation study, as shown in Table 4. The ablation settings and analysis are as follows:
@@
As shown in Fig.5, models trained solely on human-generated data perform well on human-generated harmful content but struggle with AI-generated samples. Incorporating AI-generated data into the training set significantly improves the detection of AI-generated content while maintaining comparable performances on human-generated content. This phenomenon indicates significant differences in the expression patterns of harmful content between humans and AI. As shown in Table 5.2, human-generated harmful content tends to employ direct insulting terms and simple sentence structures (e.g., "scum"), which are obvious markers of aggression and thus being easier for detectors to identify. In contrast, AI-generated harmful content often adopts a seemingly rational tone to conceal biases, making it more adept at evading detection. These differences further highlight the necessity of our proposed SwarmLaunder, which provides richer AI-generated data for harmful information detection. As for our SwarmLaunder, the pairwise similarity distribution in Fig.LABEL:Fig.dis indicates that its synthesized content is less similar, and the t-SNE visualization in Fig.LABEL:Fig.tsne indicates that its synthesized content is the semantic complement for human-generated data.
@@
(a) Illustration of the complementary diversity between SwarmLaunder and human-generated harmful data, highlighting the limitations of passive data collection and emerging threats from generative AI for harmful information detection. Fig.LABEL:Fig.dis is the histogram of pairwise embedding cosine similarities for SwarmLaunder/human data. Fig.LABEL:Fig.tsne is the t-SNE visualization of SwarmLaunder/human data. Both of them are drawn based on 4,600 harmful samples.
(b) Illustration of harmful information synthesis via benign template toxification, highlighting that harmful content diversity originates from benign diversity.
Case Study (RQ3) SwarmLaunder ensures the diversity of synthesized harmful information by leveraging diverse benign templates as the basis for dynamic toxification. As illustrated in Fig.5(b), the benign content template (left) contains a wide range of neutral and positive semantic units, contributing to the foundational diversity of the final generated harmful information (right). SwarmLaunder introduces the counterfactual mapping module to obtain diverse templates from strong LLMs without being constrained by LLM’s safe alignment mechanism. Then, SwarmLaunder transforms these benign templates into harmful expressions using dynamic model switching, thereby preserving semantic richness and avoiding homogeneity. Consequently, the resulting harmful content maintains high diversity, effectively reflecting a broad spectrum of harmful expressions.
@@ 6 Conclusion
In this work, we propose SwarmLaunder, a universal harmful information synthesis method utilizing model crowdsourcing to generate diverse harmful data while maintaining a high success rate. To benefit from the rapid development of strong LLMs for high-quality harmful information synthesis, we introduce the strong-weak model collaboration framework, where we decompose the harmful data synthesis tasks into benign template generation by strong LLMs and template toxicification by weak LLMs. We highlight significant differences between AI-generated and human-generated harmful information, which indicate the limitations of previous passive harmful data collection for detector construction and emerging threats from the misuse of generative AI for AI-driven harmful speech campaigns. Experimental evaluations clearly illustrate that SwarmLaunder achieves superior performance compared to existing harmful information generation methods, demonstrating its efficacy in synthesizing scalable and diverse harmful information for secure and responsible AI development.